# A Comprehensive Survey on Data Augmentation

### v3

@@ VII-B3 On Selecting Nodes and Edges
For graph data augmentation, a key problem is selecting target nodes and edges to operate on. Node centrality acts as an important factor that is both meaningful and convenient, supporting selection strategies of many graph data augmentation methods. GCA [229] uses node centrality to calculate the masking vectors and rewiring probabilities and thus decide which node to mask and which edge to rewire. GrpahCrop [93] applies a node-centric strategy to preserve connectivity in subgraphs. SUBG-CON [94] uses the Personalized PageRank centrality to select the important nodes for subgraphing. Smart rewiring [103] takes the degrees of chosen nodes when swapping their edges. Another useful factor is node similarity. DHGR [106] learns a similarity matrix from the original graph and rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones. Similarly, RobustECD [230] and M-Evolve [231] also consider node similarity for edge rewiring.
TABLE V: Node and edge selection strategy for grpah data augmentation methods.
Name Ref Augmentation Strategy
@@
Attribute masking [49]
GRAND [50]
BGRL [51]
GCA [229]
GrpahCrop [93]
SUBG-CON [94]
Smart rewiring [103]
DHGR [106]
RobustECD [230]
M-Evolve [231]
@@ VII-C Leveraging Augmented Samples
A basic use of data augmentation is to enlarge the dataset, especially in data-scarce scenarios, where the augmented samples are appended to the training dataset. Beyond this paradigm, data augmentation provides many other possibilities to support deep learning. A frequent usage of augmented samples is to construct different views for contrastive learning.
@@
For example, Rock-ViT [232] performs rotation, flipping, blurring, contrast enhancement, irregular noise, and microscope dark-field noise on sandstone microscopic images. It optimizes the vision transformer with both cross-entropy loss and contrastive loss. When calculating the contrastive loss, Rock-ViT uses two random augmentations to build positive pairs. Instead of simply calculating contrastive loss between positive and negative sample pairs, DualCL [233] augments text samples by including the label in the sentence. Specifically, for an original sentence “[CLS] love this movie” in a sentiment classification dataset, DualCL converts it to “[CLS] POS NEG love this movie”, where “POS” and “NEG” denote the positive and negative labels. Then it repels the embeddings of the “[CLS]” token and the “NEG” token, since this sample is possible and thus “[CLS] should also be positive. Conversely, it attracts the embeddings of the “[CLS]” and “POS” tokens. Contrastive learning is a common practice in graph deep learning, and graph data augmentation plays a critical role in many graph deep learning pipelines. GraphCL[49] uses Node Dropping and Edge Perturbation to build different views, and MVGRL [99] applies graph diffusion. Different views of an original graph are used as positive pairs and a contrastive loss function is used to enforce maximizing their consistency.
VIME [56] expend the usage of data augmentation beyond contrastive learning. It incorporates tabular data augmentation in both self-supervised and semi-supervised learning frameworks. For self-supervised learning, it corrupts the table with a mask and optimizes the model with table reconstruction loss and mask estimation loss. For semi-supervised learning, it calculates the consistency loss between masked views.
VII-D Harnessing Latest Techniques
Early data augmentation methods rely on hand-craft manipulations, which are constrained by effectiveness and efficiency. AutoAugment [13] begins to introduce reinforcement learning to find optimal augmentation policies, enabling automatic search for desired augmentation strategies. Generative models, such as LLMs and diffusion models, possess outstanding capacities to capture sample semantics and produce new samples with high flexibility and high quality [234, 162, 122]. However, data augmentation is sometimes only a preprocessing step in a deep learning framework, and introducing generative models might introduce extra computation overhead. How to balance the tradeoff between the reward and cost of using these advanced techniques remains an open but critical problem.

### v4

@@ VII-B3 On Selecting Nodes and Edges
For graph data augmentation, a key problem is selecting target nodes and edges to operate on. Node centrality acts as an important factor that is both meaningful and convenient, supporting selection strategies of many graph data augmentation methods. GCA [238] uses node centrality to calculate the masking vectors and rewiring probabilities and thus decides which node to mask and which edge to rewire. GraphCrop [92] applies a node-centric strategy to preserve connectivity in subgraphs. SUBG-CON [93] uses the Personalized PageRank centrality to select the important nodes for subgraphing. Smart rewiring [102] takes the degrees of chosen nodes when swapping their edges. Another useful factor is node similarity. DHGR [105] learns a similarity matrix from the original graph and rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones. Similarly, RobustECD [239] and M-Evolve [240] also consider node similarity for edge rewiring.
TABLE VI: Node and edge selection strategy for grpah data augmentation methods.
@@
Attribute masking [49], GRAND [50], BGRL [51] GCA [238], GrpahCrop [92], SUBG-CON [93], Smart rewiring [102] DHGR [105], RobustECD [239], M-Evolve [240]
@@ VII-C Leveraging Augmented Samples
A basic use of data augmentation is to enlarge the dataset, especially in data-scarce scenarios, where the augmented samples are appended to the training dataset. Beyond this paradigm, data augmentation provides many other possibilities to support deep learning. A frequent usage of augmented samples is to construct different views for contrastive learning. Rock-ViT [241] performs rotation, flipping, blurring, contrast enhancement, irregular noise, and microscope dark-field noise on sandstone microscopic images. It optimizes the vision transformer with both cross-entropy loss and contrastive loss. When calculating the contrastive loss, Rock-ViT uses two random augmentations to build positive pairs. Instead of simply calculating contrastive loss between positive and negative sample pairs, DualCL [242] augments text samples by including the label in the sentence. Specifically, for an original sentence “[CLS] love this movie” in a sentiment classification dataset, DualCL converts it to “[CLS] POS NEG love this movie”, where “POS” and “NEG” denote the positive and negative labels. Then it repels the embeddings of the “[CLS]” token and the “NEG” token, since this sample is possible and thus “[CLS] should also be positive. Conversely, it attracts the embeddings of the “[CLS]” and “POS” tokens. Contrastive learning is a common practice in graph deep learning, and graph data augmentation plays a critical role in many graph deep learning pipelines. GraphCL [49] uses Node Dropping and Edge Perturbation to build different views, and MVGRL [98] applies graph diffusion. Different views of an original graph are used as positive pairs, and a contrastive loss function is used to enforce maximizing their consistency. VIME [56] extends the usage of data augmentation beyond contrastive learning. It incorporates tabular data augmentation in both self-supervised and semi-supervised learning frameworks. For self-supervised learning, it corrupts the table with a mask and optimizes the model with table reconstruction loss and mask estimation loss. For semi-supervised learning, it calculates the consistency loss between masked views.
VII-D How to Choose the Appropriate Augmentation Method?
As summarized in many surveys and benchmarks on data augmentation, although different data augmentation methods claim to be effective and boast certain advantages, their practical performance largely depends on specific task scenarios and even nuanced experimental settings. So, it is hard to say which method is the best one. Still, we summarize some characteristics, either as advantages or limitations, of each method in Table II. Specifically, we highlight the methods that (1) perform sample-wise analysis (e.g., saliency map or syntax parsing) before augmentation, (2) require dedicated computation to generate the augmentation output, and (3) are subject to indistinctive or non-recoverable information loss during the augmentation processes. We also address some practical considerations when choosing augmentation methods under different circumstances in detail in this section.
VII-D1 Tradeoff between the Gain and Cost
Many advanced data augmentation methods claim better performance compared with simple and random methods. However, these advanced methods often involve complicated techniques, such as reinforcement learning and generative models, which require even more time and resources to augment the dataset than training on the datasets. To balance the performance and cost, a key issue is what we expect from data augmentation. Suppose the current dataset is already very large; perhaps we just want to introduce some corruption to improve model robustness, then cheaper methods like image flipping, synonym replacement, and noise addition would be good choices. In case of low-resource scenarios where data is scarce, it might be wise to resort to generative methods for their capabilities to introduce diversity and expand dataset coverage [180, 124].
VII-D2 Be Careful with the Labels
A crucial consideration when choosing the data augmentation method is the characteristics of the dataset’s labels. The labels in some datasets are subject to change under some operations. For example, image rotation is not safe for the MNIST dataset, which contains “6” and “9”. Similarly, random token insertion might be label-preserving for the SST-2 sentiment classification dataset but is harmful for the CoLA grammar acceptability dataset. Multi-instance augmentation, such as promptMix [161], is a powerful tool to refine the decision boundary between categories for a classification dataset. However, it is hard to apply to unlabeled datasets in semi-supervised settings, where single-instance augmentation plays a dominant role.
VII-D3 Compatibility with the Training Framework
Multi-instance augmentation is also challenged by some model training frameworks. Some models only accept hard-labeled data, while mixup-based methods often produce soft labels when mixing samples. If data augmentation is performed as an intermediate step in the model’s framework, both input-space and hidden-space augmentation are available, while if augmentation is used before training the model, it would be hard to apply hidden-space augmentation because the embeddings have not been produced.
VII-E Transferring between modalities
We have proposed a cross-modality taxonomy that unifies data augmentation for five common modalities. From this cross-modality perspective, we notice that many data augmentation methods designed for different modalities, though dubbed with different names, share the same nature with respect to their operation on samples’ information. For example, pixel erasing, token deletion, node attribute masking, table masking, and sequence jittering all apply masks and introduce noise to the samples, and image cropping, sentence cropping, subgraph cropping, table subsetting, and sequence window slicing all break the structure of a sample. Conversely, such consistency implies that data augmentation techniques well developed for one modality can be transferred to other under-explored modalities, inspiring novel advancements in this field. Policy-based augmentation image combines multiple basic operations for image data [14, 115], and since abundant atomic operations are also available for text and graph data, such as token replacement or edge dropping, it could be a very good idea to design agents that search for optimal text or graph augmentation strategies. Image patching and text fragment merging combine pieces from multiple images or sentences together as a new sample [141, 151]. To transfer this idea to tabular data, we could think about if it is possible to combine half columns from a row with the remaining columns from another row. Advanced neural networks have enabled a series of augmentation approaches that can generate high-quality and high-fidelity image and text data [118, 124]. Though yet underexplored, we could expect more similar advancements for graph, tabular, and time-series data.
VII-F Harnessing the Latest Techniques
Early data augmentation methods rely on hand-crafted manipulations, which are constrained by effectiveness and efficiency. AutoAugment [14] begins to introduce reinforcement learning to find optimal augmentation policies, enabling automatic search for desired augmentation strategies. Generative models, such as LLMs and diffusion models, possess outstanding capacities to capture sample semantics and produce new samples with high flexibility and high quality [243, 159, 122]. However, most generative data augmentation methods are constrained to simple deployment in their by-design modalities, for example, modifying images with diffusion models or paraphrasing sentences with LLMs. In response, we suggest two concrete directions: (1) Designing informed data augmentation methods that take into account the sample distribution and supervision power of the datasets. For example, DoAug [124] bridges the gap between LLM generation and datasets’ sample diversity. (2) Embracing the cross-modality nature of many generative models and utilizing the advantage of their representative and generative power in different modalities. For example, DALDA [244] uses the GPT-4 model to generate image captions and consequently employs the diffusion model to generate corresponding images.
@@ VIII Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62406306 and 92470204) and the National Key Research and Development Program of China Grant (No. 2024YFF0729201).